# DocKylin: A Large Multimodal Model for Visual Document Understanding with Efficient Visual Slimming

### v3

@@ Introduction
With the substantial growth in data and model sizes, as well as alignment with human preferences, current Large Language Models (LLMs) (OpenAI 2022; Touvron et al. 2023; The Vicuna Team 2023; Bai et al. 2023a; Team 2023; Yang et al. 2023) have demonstrated remarkable capabilities in reasoning, common sense understanding, expertise in specialized knowledge, zero/few-shot learning, and instruction following, shedding light on the potential for Artificial General Intelligence. Multimodal Large Language Models (MLLMs) aim to enhance these powerful LLMs by endowing them with multimodal capabilities, enabling the integration of diverse modalities beyond text alone.
@@
Although there has been significant progress in the development of MLLMs, their performance in visual document understanding (VDU) tasks remains suboptimal (Liu et al. 2023b; Zhang et al. 2024). A primary reason for this is their support for input resolutions of only up to $448\times 448$, which is inadequate for high-resolution, fine-grained document images. Numerous efforts (Yu et al. 2024; Liu et al. 2024b; Li et al. 2023b; Wei et al. 2023a; Bai et al. 2023b; Ye et al. 2023a; Lin et al. 2023; Wei et al. 2024) have recently been made to increase the input resolution to address this limitation and achieved notable improvements. However, these efforts still face several challenges: 1) Increasing the resolution also escalates the redundant regions within documents, which results in an increased number of redundant visual tokens, complicating the task for LLMs to identify the correct answers. 2) Some methods (Li et al. 2023b; Wei et al. 2023a, 2024; Bai et al. 2023b) employ a fixed rectangular resolution, which can lead to text distortion given the typically large aspect ratios of document images. Moreover, scaling smaller images to larger ones results in inefficiencies. 3) To manage the extended sequences of visual tokens resulting from high resolution, most methods (Li et al. 2023b; Yu et al. 2024; Liu et al. 2024b; Bai et al. 2023b) employ a fixed compression ratio or extract a predetermined number of tokens. While this approach may be suitable for natural scene images, it is less effective for document scenarios where content density varies significantly. For example, images of scientific papers are information-dense, whereas images of identification cards are comparatively sparse.
@@
To tackle these challenges, we propose DocKylin, a document-centric MLLM that incorporates several innovative designs: Firstly, we propose a parameter-free Adaptive Pixel Slimming (APS) preprocessing technique that utilizes gradient information to identify and eliminate redundant regions within document images, thereby reducing the proportion of redundant pixels and enhancing computational efficiency. Secondly, we adopt a more flexible image encoding strategy by capping the maximum number of pixels instead of fixing resolutions. This allows for variable resolutions and aspect ratios. Our lightweight Swin encoder (Liu et al. 2021) processes high-resolution document images holistically, ensuring feature coherence. Thirdly, we propose a Dynamic Token Slimming (DTS) method based on a dual-center clustering approach, which is also parameter-free. DTS filters useful tokens from a vast pool of visual tokens, obtaining visual sequences with reduced and dynamic lengths. Incorporating these innovations, we developed our DocKylin, which demonstrates superior performance across multiple VDU benchmarks. Extensive experiments demonstrate that DocKylin outperforms existing methods across various VDU benchmarks. Furthermore, both the proposed APS and DTS are parameter-free and modular, facilitating easy integration into existing MLLMs to achieve enhanced performances.
@@
Visual Document Understanding
Visual Document Understanding (VDU) (Cui et al. 2021; Zhang et al. 2023; Xu et al. 2020; Kim et al. 2022; Li et al. 2021) centers on the automated processing, interpretation, classification, and extraction of information from document images with varied typesetting formats. This task is particularly challenging due to the diverse layouts, often poor image quality, and complex structures within the documents.
Current advanced VDU methods can be broadly categorized into OCR-dependent and OCR-free approaches. OCR-dependent methods (Huang et al. 2022; Xu et al. 2021, 2020; Appalaraju et al. 2021, 2024; Tang et al. 2023; Gu et al. 2022; Wang, Jin, and Ding 2022; Li et al. 2021) leverage external Optical Character Recognition engines to extract text and layout information from documents. Such additional OCR step complicates the workflow and makes performance heavily reliant on OCR accuracy. Errors in OCR step can be accumulated to the subsequent understanding model. In contrast, OCR-free methods (Kim et al. 2022; Davis et al. 2022; Yu et al. ) process document images directly as input and produce end-to-end results without the need for an additional OCR step.
Text-centric MLLMs
Leveraging large language models (LLMs) trained on extensive datasets, emerging multimodal LLMs have recently demonstrated significant potential in visual document understanding (VDU) tasks. However, their performance still falls short of expectations (Liu et al. 2023b; Shi et al. 2023; Tang et al. 2024b). Therefore, there has been a surge of recent work studying how to enhance the perception, recognition, and understanding abilities of MLLMs in such challenging scenario. Some approaches (Tanaka et al. 2024; Kim et al. 2023) leverage additional Optical Character Recognition (OCR) engines to recognize text in images and input it into MLLMs. Others (Li et al. 2023b; Liu et al. 2024b; Bai et al. 2023b; Dong et al. 2024; Feng et al. 2023a; Ye et al. 2023a; Wei et al. 2023a) enhance the fine-grained perceptual capabilities of MLLMs by maintaining the high resolution of the input images. This includes strategies such as splitting the image into small pieces to accommodate the VIT’s requirement for fixed and small-size inputs and utilizing visual encoders that are more suitable for text images. Authors of (Feng et al. 2023b; Wang et al. 2023) explore learning text grounding for MLLMs, while (Hu et al. 2024; Tang et al. 2024a) investigate the use of more superior and larger scale instruction data to boost performance. Although these efforts have significantly improved the performance of MLLMs on text-centric tasks, there is still substantial room for improvement.
@@
Maintaining high-resolution input images is a straightforward and effective method to enhance the performance of MLLMs on text images. However, increased resolutions inevitably lead to a longer visual sequence. There are various methods proposed for compressing the length of visual sequences in current MLLMs. One of the earlier and commonly used approaches (Dai et al. 2024; Li et al. 2023a; Alayrac et al. 2022) involves utilizing a group of learnable query tokens to extract information in a cross-attention manner. Another method (Chen et al. 2023a; Yu et al. 2024) involves concatenating adjacent tokens along the channel dimension to reduce sequence length. Additionally, convolutional neural networks (CNNs) have also been proven effective for visual token compression (Cha et al. 2023; Hu et al. 2024). However, all these methods either extract tokens of a fixed length or compress them at a fixed ratio, which is unsuited for document images, where the content density can vary significantly. A token sequence that is too short or a compression ratio that is too high can result in an inadequate representation of some document images. Conversely, longer token sequences or lower compression rates can reduce computational efficiency.
Recent studies (Cai et al. 2024) have shown that the optimal length of visual token sequences varies for each sample. Consequently, some recent approaches have explored the adaptive compression of visual sequences. Most of these methods (Liu et al. 2024a; Shang et al. 2024; Chen et al. 2024; Lin et al. 2024) dynamically discard nonessential tokens using attention scores between visual tokens or between visual tokens and class tokens. However, considering the differences between text and natural objects with complete semantics, this paradigm may not be suitable for text images. For text-centric images, HRVDA (Wei et al. 2024) employs text position annotations to train a separate visual token filtering network, which requires introducing additional parameters and training procedures. Compared to the methods above, we explore alternative dynamic compression schemes specifically for document images, which include pixel-level redundancy removal and unsupervised clustering for filtering out nonessential tokens.
@@
As illustrated in Fig. 1, the architecture of DocKylin incorporates an image encoder, an MLP layer that aligns the visual modality to the language modality, and a LLM. Building on this framework, DocKylin introduces the following enhancements: 1) An Adaptive Pixel Slimming (APS) module preprocesses images before they are fed into the image encoder, which can removes redundant regions and reduces resolution while increasing the proportion of informative pixels. 2) a more lightweight and flexible image encoder by capping the maximum number of pixels instead of fixing resolutions. 3) A Dynamic Token Slimming (DTS) module handles the visual tokens produced by the visual encoder and MLP layer. DTS further removes nonessential tokens to enhance both performance and efficiency.
@@
Perceiving textual information in images typically requires maintaining a high resolution of input, especially for text-rich document images. This poses significant challenges for current resource-intensive MLLMs. To address this issue, existing methods employ strategies such as dividing the images into patches for processing (Liu et al. 2024b; Ye et al. 2023a), employing visual encoders that support high resolution (Wei et al. 2023a), leveraging OCR results as additional inputs (Tanaka et al. 2024; Kim et al. 2023), or applying compression transformations to the images beforehand (Feng et al. 2023a).
@@
Figure 2: Schematic representation of the Adaptive Pixel Slimming module. It effectively reduces the resolution of document images by removing redundant regions.
Here, we propose a novel approach called Adaptive Pixel Slimming (APS) to efficiently reduce the resolution of document images, which is orthogonal to the approaches mentioned above. APS is based on the observation that, as illustrated in Fig. 2, while document images usually exhibit high resolution, they commonly contain a significant proportion of redundant regions. These regions not only increase the computational burden on the visual encoder but may also pose challenges (Liu et al. 2024b; Wei et al. 2024) to the reasoning capabilities of the LLMs by the increased redundant visual tokens. APS is thus designed to identify and remove these regions. Specifically, as shown in Fig. 2, during the gradient extraction stage, we utilize the Sobel operator to extract the gradient map $I_{G}$ from the original image $I_{S}$. Regions with low gradient values are considered smooth and devoid of significant information. Therefore, in the redundant region determination stage, we identify the redundant regions in the horizontal and vertical directions by locating areas with consistently low gradient values. For instance, to identify redundant regions in the image’s horizontal direction: if the sum of gradient values for a specific row is below a predefined threshold, we classify that row as redundant. A sequence of contiguous redundant rows constitutes a redundant region in the horizontal direction. The computation for the vertical direction follows a similar procedure. The final image can be obtained by removing the identified redundant rows and columns. For more details on the algorithm and parameter settings, please refer to the supplementary materials.
@@
Most existing MLLMs (Li et al. 2023b) utilize a fixed rectangular resolution, such as 224$\times$224 or 336$\times$336. This approach distorts the original aspect ratio of the images, leading to text deformation and inefficiencies when upscaling images smaller than the preset resolution. While some methods (Ye et al. 2023a; Yu et al. 2024) allow for more flexible aspect ratios and resolutions, they often employ a patch-based processing strategy. Such a strategy may be suitable for natural scene images where the continuity and context of visual elements are less critical. However, it is less effective for document images (Liu et al. 2024b) that contain dense text. Segmenting such images can disrupt the textual content, severing lines of text or splitting words and sentences across different patches. This discontinuity can impair the ability to accurately recognize and interpret textual information essential for effective document image understanding.
@@
Figure 3: Schematic diagram of the proposed Similarity Weighted Aggregation module. The aggregation module aggregates nonessential tokens into essential tokens through similarity weighted summation.
Although ViT pre-trained with CLIP (Radford et al. 2021) demonstrates impressive generalization capabilities for natural scene images, they perform suboptimally on dense document images (Li et al. 2024). Furthermore, their self-attention mechanism, which has a computational complexity that scales quadratically with image resolution, makes it challenging to support high-resolution inputs. Consequently, we opt to use the more efficient Swin Transformer (Liu et al. 2021) as our visual encoder, which has been pre-trained on full-text recognition tasks (Kim et al. 2022). It enables us to process an entire image as a single input.
@@
11:num1 =num_of_same_tokens(top-R_similar_tokens, cluster1)
12:num2 =num_of_same_tokens(top-R_similar_tokens, cluster2)
@@
Dynamic Token Slimming
Although the APS module has been effective in removing some redundant regions, it is limited to processing entire rows or columns and is ineffective for more complex scenarios. In contrast, our Dynamic Token Slimming (DTS) method offers a more flexible approach to eliminate redundant visual content at the token level. The design of the DTS is based on the premise that a well-trained visual encoder should be capable of effectively distinguishing between essential and nonessential regions within an image, with the features corresponding to different regions exhibiting sufficient discriminative properties. We propose to separate these two types of tokens directly through clustering and then aggregate nonessential tokens into essential tokens. We next detail the proposed DTS, which comprises two main processes: Dual-center K-Means Clustering and Similarity Weighted Aggregation.
Dual-center K-Means Clustering. To distinguish tokens into essential and nonessential types without introducing additional parameters, we apply K-Means clustering (Lloyd 1982) with two centers to the output sequence from the visual encoder, as illustrated in Fig 1. However, after obtaining the two clusters, it is unclear which cluster contains the essential tokens and which contains the non-essential tokens. Drawing on findings from (Liu et al. 2024b), nonessential tokens typically lack uniqueness and are often similar to other tokens. Therefore, we identify potential nonessential tokens by evaluating the similarity of each visual token with all other ones. The cluster with fewer of these nonessential tokens is determined to be the one containing the essential tokens. For more implementation details, please refer to Algorithm 1.
Table 1: Quantitative accuracy (%) comparison of DocKylin with existing general-purpose MLLMs on document image benchmarks. Accuracy proposed in (Zhang et al. 2024) is adopted as the metric for all benchmarks.
@@
mPLUG-Owl (Ye et al. 2023b) CLIP-L (0.3B) LLaMA-7B 7.4 20.0 7.9 0.2 0.1 0.3
@@
LLaVA1.5 (Liu et al. 2023a) CLIP-L (0.3B) LLaMA-7B 8.5 14.7 9.3 0.2 1.7 2.5
@@
Monkey (Li et al. 2023b) CLIP-G (2B) Qwen-7B 50.1 25.8 54.0 24.1 41.9 19.9
@@
mPLUG-Doc (Ye et al. 2023b) CLIP-L (0.3B) LLaMA-7B 62.2 38.2 57.4 188.8 42.6 26.9
UReader (Ye et al. 2023a) CLIP-L (0.3B) LLaMA-7B 65.4 42.2 65.7 221.7 49.5 29.4
@@
HRVDA (Wei et al. 2024) Swin LLaMA2-7B 72.1 43.5 67.6 211.5 63.2 31.2
Vary (Wei et al. 2023a) ViTDet-B & CLIP-L (0.4B) Vicuna-7B 76.3 - 66.1 - - -
TextMonkey (Liu et al. 2024b) CLIP-G (2B) Qwen-7B 71.5 - 65.5 - 61.6 30.6
@@
LayoutLLM (Huang et al. 2022) LayoutLMv3-L (0.3B) LLaMA2-7B 74.3 - - - - -
@@
Similarity Weighted Aggregation. After obtaining essential and nonessential tokens, how to handle the nonessential tokens becomes a crucial issue. Inspired by (Kong et al. 2022; Liang et al. 2021; Wei et al. 2023b), to mitigate potential clustering errors and prevent information loss, we do not discard the nonessential tokens outright. Instead, we aggregate them into the essential tokens to minimize the potential loss of effective information while enhancing the model’s processing efficiency. As illustrated in Fig. 3, for each nonessential token from the clustering step, we find the most similar essential token. Each essential token may correspond to zero or multiple nonessential tokens. If an essential token has corresponding nonessential tokens, we perform a similarity weighted summation of these tokens to obtain an aggregated token.
@@
The pseudo-code for the specific implementation can be found in the supplementary materials.
Figure 4: Visualization from DocKylin. The redundant pixels and nonessential tokens identified by Adaptive Pixel Slimming and Dynamic Token Slimming are highlighted in light blue.
Table 3: Comparisons with current SOTA parameter-free visual token compression methods. Accuracy proposed in (Zhang et al. 2024) is adopted as the metric for all benchmarks.
@@
FastV (K=2, R=50%) (Chen et al. 2024) 306 8.4 306 14.4 306 2.0
@@
APS+DTS (ours) 298 9.7 270 14.7 296 3.1
@@
During the pre-training phase, the model was trained for 450,000 iterations with a batch size of 8. The learning rate decayed from 1e-4 to 1e-5 using cosine annealing, which took approximately 40 A800 GPU-days. In the instruction tuning phase, the model underwent 300,000 iterations with the same batch size and learning rate as in the pre-training phase. This phase further consumed approximately 24 A800 GPU-days. Our Adaptive Pixel Slimming method is employed in both phases, while the Dynamic Token Slimming (DTS) is used only during the instruction tuning phase. The maximum allowed number of pixels, MAX_SIZE, is set to 1728$\times$1728. We initialize our visual encoder with Donut-Swin (0.07B) (Kim et al. 2022), pre-trained on the reading texts task, and use Qwen-7B-Chat (Bai et al. 2023a) as our LLM.
@@
We first compare our model with several existing general-purpose multimodal large language models (MLLMs). As these models sometimes tend to provide detailed responses, following (Zhang et al. 2024), we adopt accuracy as our metric: a response is considered correct if it contains the complete answer. The results, as shown in Table 1, demonstrate that DocKylin achieves a significant advantage over the competing models. This highlights that VDU tasks remain challenging for current general-purpose MLLMs. Note that descriptions of all these benchmarks are included in the supplementary materials.

### v4

@@ Introduction
With the substantial growth in data and model sizes, as well as alignment with human preferences, current Large Language Models (LLMs) (OpenAI 2022; Touvron et al. 2023; The Vicuna Team 2023; Bai et al. 2023a; The InternLM Team 2023; Yang et al. 2023) have demonstrated remarkable capabilities in reasoning, common sense understanding, expertise in specialized knowledge, zero/few-shot learning, and instruction following, shedding light on the potential for Artificial General Intelligence. Multimodal Large Language Models (MLLMs) aim to enhance these powerful LLMs by endowing them with multimodal capabilities, enabling the integration of diverse modalities beyond text alone (Yin et al. 2024; Fu et al. 2024).
@@
Although there has been significant progress in the development of MLLMs, their performance in visual document understanding (VDU) tasks remains suboptimal (Liu et al. 2023a; Zhang et al. 2024b). A primary reason for this is their support for input resolutions of only up to $448\times 448$, which is inadequate for high-resolution, fine-grained document images. Numerous efforts (Yu et al. 2024; Liu et al. 2024d; Li et al. 2024b; Wei et al. 2025; Bai et al. 2023b; Ye et al. ; Lin et al. 2023; Liu et al. 2024a) have recently been made to increase the input resolution to address this limitation and achieved notable improvements. However, these efforts still face several challenges: 1) Increasing the resolution also escalates the redundant regions within documents, which results in an increased number of redundant visual tokens, complicating the task for LLMs to identify the correct answers. 2) Some methods (Li et al. 2024b; Wei et al. 2025; Liu et al. 2024a; Bai et al. 2023b) employ a fixed rectangular resolution, which can lead to text distortion given the typically large aspect ratios of document images. Moreover, scaling smaller images to larger ones results in inefficiencies. 3) To manage the extended sequences of visual tokens resulting from high resolution, most methods (Li et al. 2024b; Yu et al. 2024; Liu et al. 2024d; Bai et al. 2023b) employ a fixed compression ratio or extract a predetermined number of tokens. While this approach may be suitable for natural scene images, it is less effective for document scenarios where content density varies significantly. For example, images of scientific papers are information-dense, whereas images of identification cards are comparatively sparse.
@@
To tackle these challenges, we propose DocKylin, a document-centric MLLM that incorporates several innovative designs: Firstly, we propose a parameter-free Adaptive Pixel Slimming (APS) preprocessing technique that utilizes gradient information to identify and eliminate redundant regions within document images, thereby reducing the proportion of redundant pixels and enhancing computational efficiency. Secondly, we adopt a more flexible image encoding strategy by capping the maximum number of pixels instead of fixing resolutions. This allows for variable resolutions and aspect ratios. Our lightweight Swin encoder (Liu et al. 2021) processes high-resolution document images holistically, ensuring feature coherence. Furthermore, we introduce a parameter-free Dynamic Token Slimming (DTS) method based on dual-center clustering. DTS efficiently filters informative tokens from a large set of visual tokens, resulting in reduced and dynamically sized visual sequences. Incorporating these innovations, we developed our DocKylin, which demonstrates superior performance across multiple VDU benchmarks. Extensive experiments demonstrate that DocKylin outperforms existing methods across various VDU benchmarks. Furthermore, both the proposed APS and DTS are parameter-free and modular, facilitating easy integration into existing MLLMs to achieve enhanced performances.
@@
Text-centric MLLMs for VDU
Visual document understanding (VDU) (Cui et al. 2021; Zhang et al. 2023a; Xu et al. 2020; Kim et al. 2022) centers on the automated processing, interpretation, classification, and extraction of information from document images with varied typesetting formats. This task is particularly challenging due to the diverse layouts (Cheng et al. 2023), often poor image quality (Zhang et al. 2024a, 2023b, 2022), and complex structures within the documents.
Leveraging large language models (LLMs) trained on extensive datasets, emerging MLLMs have recently demonstrated significant potential in VDU tasks. However, their performance still falls short of expectations (Liu et al. 2023a; Zhang et al. 2024b; Tang et al. 2024b; Shan et al. 2024). Therefore, there has been a surge of recent work studying how to enhance the perception, recognition, and understanding abilities of MLLMs in such challenging scenarios.
Some approaches (Tanaka et al. 2024; Kim et al. 2023; Lu et al. 2024; Liao et al. 2024) leverage additional OCR engines to recognize text in images and input it into MLLMs. Others (Li et al. 2024b; Liu et al. 2024d; Bai et al. 2023b; Dong et al. 2024; Feng et al. 2023a; Ye et al. ; Wei et al. 2025; Zhao et al. 2024a, b) enhance the fine-grained perceptual capabilities of MLLMs by maintaining the high resolution of the input images. This includes strategies such as splitting the image into small pieces to accommodate the ViT’s requirement for fixed and small-size inputs and utilizing visual encoders that are more suitable for text images. Authors of (Feng et al. 2023b; Wang et al. 2023) explore learning text grounding for MLLMs, while (Hu et al. 2024; Tang et al. 2024a) investigate the use of more superior and larger scale instruction data to boost performance. Although these efforts have significantly improved the performance of MLLMs on text-centric tasks, there is still substantial room for improvement.
@@
Maintaining high-resolution inputs is a straightforward and effective method to enhance the performance of MLLMs on text images. However, increased resolutions inevitably lead to a longer visual sequence. To eliminate this problem, various methods are proposed for compressing the length of visual sequences in current MLLMs. One of the earlier and commonly used approaches (Li et al. 2023; Wang et al. 2024) involves utilizing a group of learnable query tokens to extract information in a cross-attention manner. Another method (Chen et al. 2023a; Yu et al. 2024) involves concatenating adjacent tokens along the channel dimension to reduce sequence length. Additionally, convolutional neural networks have also been proven effective for visual token compression (Cha et al. 2024; Hu et al. 2024). However, all these methods either extract tokens of a fixed length or compress them at a fixed ratio, which is unsuited for document images, where the content density can vary significantly. A token sequence that is too short or a compression ratio that is too high can result in an inadequate representation. Conversely, longer token sequences or lower compression rates can reduce computational efficiency.
Recent studies (Cai et al. 2024) have shown that the optimal length of visual token sequences varies for each sample. Consequently, some recent approaches have explored the adaptive compression of visual sequences. Most of these methods (Liu et al. 2024c; Shang et al. 2024; Chen et al. 2025; Lin et al. 2024) dynamically discard nonessential tokens using attention scores between visual tokens or between visual tokens and class tokens. However, considering the differences between text and natural objects with complete semantics, this paradigm may not be suitable for text images. For text-centric images, HRVDA (Liu et al. 2024a) employs text position annotations to train a separate visual token filtering network, which requires introducing additional parameters and training procedures. Compared to the methods above, we explore alternative dynamic compression schemes specifically for document images, which include pixel-level redundancy removal and unsupervised clustering for filtering out nonessential tokens.
@@
As illustrated in Fig. 1, the architecture of DocKylin incorporates an image encoder, an MLP layer that aligns the visual modality to the language modality, and a LLM. Building on this framework, DocKylin introduces the following enhancements: 1) An Adaptive Pixel Slimming (APS) module preprocesses images before they are fed into the image encoder, which can remove redundant regions and reduce resolution while increasing the proportion of informative pixels. 2) A more lightweight and flexible image encoder by capping the maximum number of pixels instead of fixing resolutions. 3) A Dynamic Token Slimming (DTS) module handles the visual tokens produced by the visual encoder and MLP layer. DTS further removes nonessential tokens to enhance both performance and efficiency.
In this paper, terms like ‘redundant’ and ‘nonessential’ refer to areas that can be masked without altering the document’s overall meaning, allowing both humans and models to answer questions correctly. Examples include the background of the document or color blocks in bar charts.
@@
Perceiving textual information in images typically requires maintaining a high input resolution, especially for text-rich document images. This poses significant challenges for current resource-intensive MLLMs. To address this issue, existing methods employ strategies such as dividing the images into patches for processing (Ye et al. ), employing visual encoders that support high resolution (Wei et al. 2025), leveraging OCR results as additional inputs (Tanaka et al. 2024; Kim et al. 2023), or applying compression transformations to the images beforehand (Feng et al. 2023a).
@@
Figure 2: The proposed Adaptive Pixel Slimming module. It effectively reduces the resolution of document images by removing redundant regions.
Here, orthogonal to the approaches mentioned above, we propose a novel approach called Adaptive Pixel Slimming (APS) to efficiently reduce the resolution of document images. APS is based on the observation that, as illustrated in Fig. 2, while document images usually exhibit high resolution, they commonly contain a significant proportion of redundant regions. These regions not only increase the computational burden on the visual encoder but may also pose challenges (Liu et al. 2024d, a) to the reasoning capabilities of the LLMs by the increased redundant visual tokens. APS is thus designed to identify and remove these regions. Specifically, as shown in Fig. 2, during the gradient extraction stage, we utilize the Sobel operator to extract the gradient map $I_{G}$ from the original image $I_{S}$. Regions with low gradient values are considered smooth and devoid of significant information. Therefore, in the redundant region determination stage, we identify the redundant regions in the horizontal and vertical directions by locating areas with consistently low gradient values. For instance, to identify redundant regions in the image’s horizontal direction: if the sum of gradient values for a specific row is below a predefined threshold, we classify that row as redundant. A sequence of contiguous redundant rows constitutes a redundant region in the horizontal direction. The vertical direction follows a similar procedure. The final image can be obtained by removing the identified redundant rows and columns.
@@
Most existing MLLMs (Li et al. 2024b) utilize a fixed rectangular resolution, such as 224$\times$224 or 336$\times$336. This approach distorts the original aspect ratio of the images, leading to text deformation and inefficiencies when upscaling images smaller than the preset resolution. While some methods (Ye et al. ; Yu et al. 2024) allow for more flexible aspect ratios and resolutions, they often employ a patch-based processing strategy. Such a strategy may be suitable for natural scene images where the continuity and context of visual elements are less critical. However, it is less effective for document images that contain dense text (Liu et al. 2024d; Huang et al. 2024). Segmenting such images can disrupt the textual content, severing lines of text or splitting words and sentences across different patches. This discontinuity can impair the ability to accurately recognize and interpret textual information essential for effective document image understanding.
@@
Although ViT pre-trained with CLIP (Radford et al. 2021) demonstrates impressive generalization capabilities for natural scene images, they perform suboptimally on dense text images (Li et al. 2024a; Shan et al. 2024). Furthermore, their self-attention mechanism, with its quadratic computational complexity scaling with image resolution, presents a significant challenge for processing high-resolution inputs. Consequently, we opt to use the more efficient Swin Transformer (Liu et al. 2021) as our visual encoder, which has been pre-trained on full-text recognition tasks (Kim et al. 2022). It enables us to process an entire image as a single input.
Dynamic Token Slimming
Although the APS module has been effective in removing some redundant regions, it is limited to processing entire rows or columns and is ineffective for more complex scenarios. In contrast, our Dynamic Token Slimming (DTS) method offers a more flexible approach to eliminate redundant visual content at the token level. The design of the DTS is based on the premise that a well-trained visual encoder should be capable of effectively distinguishing between essential and nonessential regions within an image, with the features corresponding to different regions exhibiting sufficient discriminative properties. We propose to separate these two types of tokens directly through clustering and then aggregate nonessential tokens into essential ones. We next detail the proposed DTS, which comprises two main processes: Dual-center K-Means Clustering and Similarity Weighted Aggregation.
Dual-center K-Means Clustering. To distinguish tokens into essential and nonessential types without introducing additional parameters, we apply K-Means clustering with two centers to the output sequence from the visual encoder, as illustrated in Fig 1. However, after obtaining the two clusters, it is unclear which cluster contains the essential tokens and which contains the non-essential tokens. Drawing on findings from (Liu et al. 2024d), nonessential tokens typically lack uniqueness and are often similar to other tokens. Therefore, we identify potential nonessential tokens by evaluating the similarity of each visual token with all other ones. The cluster with fewer of these nonessential tokens is determined to be the one containing the essential tokens. For more implementation details, please refer to Algorithm 1.
@@
11:num1 =The number of common tokens between top-R_similar_tokens and cluster1
12:num2 =The number of common tokens between top-R_similar_tokens and cluster2
@@
Figure 3: The proposed Similarity Weighted Aggregation module. It aggregates nonessential tokens into essential ones through similarity-weighted summation.
@@
mPLUG-Owl (Ye et al. 2023) CLIP-L (0.3B) LLaMA-7B 7.4 20.0 7.9 0.2 0.1 0.3
@@
LLaVA1.5 (Liu et al. 2024b) CLIP-L (0.3B) LLaMA-7B 8.5 14.7 9.3 0.2 1.7 2.5
@@
Monkey (Li et al. 2024b) CLIP-G (2B) Qwen-7B 50.1 25.8 54.0 24.1 41.9 19.9
@@
Table 1: Quantitative comparison of DocKylin with existing general-purpose MLLMs on document image benchmarks. Accuracy (%) proposed in (Zhang et al. 2024b) is adopted as the metric for all benchmarks.
@@
mPLUG-Doc (Ye et al. 2023) CLIP-L (0.3B) LLaMA-7B 62.2 38.2 57.4 188.8 42.6 26.9
UReader (Ye et al. ) CLIP-L (0.3B) LLaMA-7B 65.4 42.2 65.7 221.7 49.5 29.4
@@
HRVDA (Liu et al. 2024a) Swin LLaMA2-7B 72.1 43.5 67.6 211.5 63.2 31.2
Vary (Wei et al. 2025) ViTDet-B & CLIP-L (0.4B) Vicuna-7B 76.3 - 66.1 - - -
TextMonkey (Liu et al. 2024d) CLIP-G (2B) Qwen-7B 71.5 - 65.5 - 61.6 30.6
@@
LayoutLLM (Luo et al. 2024) LayoutLMv3-L (0.3B) LLaMA2-7B 74.3 - - - - -
@@
Similarity Weighted Aggregation. After obtaining essential and nonessential tokens, how to handle the nonessential tokens becomes a crucial issue. Inspired by (Kong et al. 2022; Liang et al. 2021; Wei et al. 2023), to mitigate potential clustering errors and prevent information loss, we do not discard the nonessential tokens outright. Instead, we aggregate them into the essential tokens to minimize the potential loss of effective information while enhancing the model’s processing efficiency. As illustrated in Fig. 3, for each nonessential token from the clustering step, we find the most similar essential token. Each essential token may correspond to zero or multiple nonessential tokens. If an essential token has corresponding nonessential tokens, we perform a similarity-weighted summation of these tokens to obtain an aggregated token.
@@
Figure 4: Visualization from DocKylin. The redundant pixels and nonessential tokens identified by Adaptive Pixel Slimming and Dynamic Token Slimming are highlighted in light blue. Zoom in for best view.
@@
MaxPool1D 288 6.7 288 13.8 288 0.5
AvgPool1D 288 7.3 288 14.5 288 1.1
FastV (K=2, R=50%) (Chen et al. 2025) 306 8.4 306 14.4 306 2.0
@@
APS+DTS w/o Aggregation (ours) 298 9.9 270 14.2 296 3.4
APS+DTS (ours) 298 9.9 270 14.9 296 3.3
Table 3: Comparisons with current SOTA parameter-free visual token compression methods. Accuracy proposed in (Zhang et al. 2024b) is adopted as the metric for all benchmarks. ‘DTS w/o Aggregation’ refers to skipping the Similarity Weighted Aggregation process mentioned earlier and directly discarding the nonessential tokens.
@@
During the pre-training phase, the model was trained for 450,000 iterations with a batch size of 8. The learning rate decayed from 1e-4 to 1e-5 using cosine annealing, which took approximately 40 A800 GPU days. In the instruction tuning phase, the model underwent 300,000 iterations with the same batch size and learning rate as in the pre-training phase. This phase further consumed approximately 24 A800 GPU days. Our Adaptive Pixel Slimming method is employed in both phases, while Dynamic Token Slimming (DTS) is used only during the instruction tuning phase. The maximum allowed number of pixels, MAX_SIZE, is set to 1728$\times$1728. We initialize our visual encoder with Donut-Swin (0.07B) (Kim et al. 2022), pre-trained on the reading texts task, and use Qwen-7B-Chat (Bai et al. 2023a) as our LLM.
@@
We first compare our model with several existing general-purpose multimodal large language models (MLLMs). As these models sometimes tend to provide detailed responses, following (Zhang et al. 2024b), we adopt accuracy as our metric: a response is considered correct if it contains the complete answer. The results, as shown in Table 1, demonstrate that DocKylin achieves a significant advantage over the competing models. This highlights that VDU tasks remain challenging for current general-purpose MLLMs. Note that descriptions of all these benchmarks are included in the supplementary materials.